OpenStudy (anonymous):

ln e^(3-4x) hellppppp

jimthompson5910 (jim_thompson5910):

same idea as last time ln(e^x) = x so... ln(e^(3-4x)) = 3 - 4x

OpenStudy (anonymous):

okay, now i see lol i didnt recognize it was the same problem..

jimthompson5910 (jim_thompson5910):

yeah just keep in mind or write it down that e^(ln(x)) = x and ln(e^x) = x

OpenStudy (anonymous):

\[3^{x+4}=7^{x-1}\]

OpenStudy (anonymous):

i have a lot of problem to solve like this i just needed help figuring out how to solve them

jimthompson5910 (jim_thompson5910):

you take the natural log of both sides, then you use the rule: ln(x^y) = y*ln(x) \[\large 3^{x+4}=7^{x-1}\] \[\large \ln\left(3^{x+4}\right)=\ln\left(7^{x-1}\right)\] \[\large (x+4)\ln\left(3\right)=(x-1)\ln\left(7\right)\]